OpenStudy (anonymous):

What polynomial has roots of -5, 2, and 4 ?

OpenStudy (cwrw238):

(x + 5)((x - 2)(x - 4) expand to get you answer

OpenStudy (anonymous):

what does that mean?

OpenStudy (cwrw238):

heres a simple example : (x - 2)(x + 1) = x(x + 1) - 2(x + 1) = x^2 + x - 2x -2 = x^2 - x - 2 then if theres another binomial (as in your example) repeat the process again this process is called the application of the distributive law

OpenStudy (amistre64):

does it have the stated roots ONLY? or can the stated roots be a subset of the roots set?

OpenStudy (amistre64):

OR, does it give you a set of options that its asking you which one has those as roots

OpenStudy (amistre64):

(x + 5)(x - 2)(x - 4) is one poly a (x + 5)(x - 2)(x - 4) is a family of polys a (x + 5)^m (x - 2)^n (x - 4)^p are even more polys a (x + 5)^m (x - 2)^n (x - 4)^p .... (x-r1) (x-r2) .... (x-rn) are even more polys

OpenStudy (anonymous):

x3 - x2 - 22x + 40 x3 + x2 - 22x - 40 x3 + 3x2 - 18x - 40 x3 - 3x2 - 18x + 40 my options

OpenStudy (amistre64):

a simple way, if factoring isnt your strong point, is to just plug in the value of the given roots to see which one makes zeros

OpenStudy (anonymous):

ok